# Learnability Window in Gated Recurrent Neural Networks

### 20 Mar 2026 v7

@@ Abstract
We develop a statistical theory of temporal learnability in recurrent neural networks, quantifying the maximal temporal horizon $\mathcal{H}_{N}$ over which gradient-based learning can recover lag-dependent structure at finite sample size $N$. The theory is built on the effective learning rate envelope $f(\ell)$, a functional that captures how gating mechanisms and adaptive optimizers jointly shape the coupling between state-space transport and parameter updates during Backpropagation Through Time. Under heavy-tailed ($\alpha$-stable) gradient noise, where empirical averages concentrate at rate $N^{-1/\kappa_{\alpha}}$ with $\kappa_{\alpha}=\alpha/(\alpha-1)$, the interplay between envelope decay and statistical concentration yields explicit scaling laws for the growth of $\mathcal{H}_{N}$: logarithmic, polynomial, and exponential temporal learning regimes emerge according to the decay law of $f(\ell)$. These results identify the envelope decay geometry as the key determinant of temporal learnability: slower attenuation of $f(\ell)$ enlarges the learnability window $\mathcal{H}_{N}$, while heavy-tailed gradient noise compresses temporal horizons by weakening statistical concentration. Experiments across multiple gated architectures and optimizers corroborate these structural predictions.
@@ 1 Introduction
Recurrent neural networks (RNNs) are fundamental models for processing sequential data, yet their ability to learn long-range temporal dependencies remains only partially understood. Gated architectures such as the LSTM and GRU have dramatically improved numerical stability and empirical performance, but it is still unclear which temporal dependencies are statistically recoverable under finite data. Most existing analyses focus on dynamical stability, spectral properties of Jacobian products, or mean-field approximations. These approaches characterize when gradients do not explode or vanish, but they do not provide a statistical criterion for when transported gradient signals remain distinguishable from noise during training. In practice, even stable gradients may be too attenuated or too noisy to carry usable information. Thus, stability alone does not determine temporal learnability.
@@
Our previous work showed that gating induces heterogeneous time scales that shape both state evolution and gradient transport [38]. Here we move beyond purely dynamical analyses and develop a statistical theory of finite-horizon learnability. The central objects are the effective learning rates $\mu_{t,\ell}$, which quantify how Backpropagation Through Time (BPTT) reweights gradient signals across temporal lags for each neuron. These quantities arise from the interaction between gate-induced transport of gradients through recurrent dynamics and the base learning rates imposed by the optimizer. Their aggregate magnitude defines the effective learning rate envelope $f(\ell)=\|\mu_{t,\ell}\|_{1}$, which governs the effective strength of lagged gradient contributions to parameter updates during training. Although effective learning rates were originally [38] derived under SGD with a fixed global learning rate $\mu$, we generalize them here to adaptive optimizers such as Adam, where each parameter receives its own time-varying rate. The generalization replaces the global $\mu$ with a neuron-specific adaptive base rate $\Lambda^{(q)}_{r,\ell}$, obtained by projecting the parameter-space preconditioner onto each neuron via a Rayleigh-quotient construction, while leaving the gate-induced transport factors unchanged.
@@
The central question addressed in this paper is: given a finite number of training sequences, up to which temporal horizon can dependencies be statistically detected? To answer this, we introduce the learnability window $\mathcal{H}_{N}$, defined as the largest lag for which gradient information remains recoverable at sample size $N$. Under heavy-tailed ($\alpha$-stable) gradient noise, empirical averages concentrate at rate $N^{-1/\kappa_{\alpha}}$ with $\kappa_{\alpha}=\alpha/(\alpha-1)$. We show that the interaction between this concentration rate and the decay geometry of the envelope $f(\ell)$ determines the scaling behavior of $\mathcal{H}_{N}$. In particular, exponential, polynomial, and logarithmic attenuation of $f(\ell)$ induce qualitatively distinct growth laws for the learnability window, providing a universal classification of temporal learning regimes.
@@
Although the detectability relation analytically separates two factors, the decay geometry of the envelope and the statistical concentration of gradient noise, these mechanisms are not independent during training. The temporal transport of gradients is shaped jointly by gating dynamics and optimizer adaptation, while the statistics of gradient fluctuations evolve with the optimization process. Empirically, regimes exhibiting slow envelope decay often co-occur with persistent heavy-tailed gradient fluctuations, reflecting the coupled evolution of state and parameter dynamics.
@@
Importantly, heavy-tailed noise does not enhance learnability. Instead, it weakens statistical concentration and therefore raises the data requirements for reliable detection of long-range dependencies. This interplay is reflected in the relation $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, which links envelope decay to the sample complexity of temporal credit assignment. Temporal learnability is thus determined by the balance between envelope attenuation and noise concentration: heterogeneous time-scale structure tends to slow the decay of $f(\ell)$ and extend the learnability window, whereas heavy-tailed gradient noise counteracts this effect by reducing statistical efficiency.
@@
Relative to prior dynamical analyses of gated RNNs, this paper makes four contributions. (i) We formalize the learnability window $\mathcal{H}_{N}$ as a finite-sample measure of recoverable temporal dependencies and incorporate heavy-tailed ($\alpha$-stable) gradient noise into the analysis of recurrent training. (ii) We derive explicit scaling laws that classify temporal learning regimes into canonical classes, linking envelope decay $f(\ell)$, sample-complexity requirements $N(\ell)$, and the growth of the learnability window $\mathcal{H}_{N}$. (iii) We generalize the effective learning rate framework from SGD to adaptive optimizers by introducing a neuron–lag projection of the parameter-space preconditioner. (iv) We empirically validate these structural predictions across multiple gated architectures.
@@ 2 Related work
The introduction of gating mechanisms in LSTM [20] and GRU [9] was pivotal for controlling temporal credit assignment in recurrent neural networks. Early theoretical works linked the forget gate to an interpretable exponential decay and controllable memory retention [15, 51], while more recent studies connect gating to continuous-time or implicit ODE perspectives [5, 16]. These contributions illuminate how gates shape state dynamics, but do not analyze how gating multiplicative structures affect parameter updates and the capacity for learning long-range dependencies. Our framework bridges that gap by deriving per-lag effective learning rates from products of Jacobians and linking them to finite-horizon learnability.
@@
The classical vanishing/exploding gradients problem motivated many stabilization techniques: clipping, spectral regularization, and orthogonality or unitary RNNs [3, 43, 1, 54, 26]. Hierarchical, dilated, or skip architectures shorten propagation paths [30, 6, 10], while continuous-time and neural ODE approaches enhance stability and expressivity [45, 18]. Works on orthogonal / dynamical isometry control the conditioning of gradient transport [7, 44] but do not address whether transported gradients still carry statistically reliable information over long lags. In contrast, we analyze statistical detectability, showing that even when products of Jacobians remain numerically stable, learning can be ineffective if the envelope decays too rapidly. Related difficulties have long been observed in tasks with delayed supervision or rewards [3], where the learning signal must be propagated across long temporal gaps; however, existing analyses typically focus on optimization stability rather than statistical detectability across lags.
@@
A related but distinct line of work concerns memory capacity in reservoir computing and echo state networks, where the recurrent dynamics are kept fixed and only a readout layer is trained [53, 24, 13]. In this setting, classical results characterize the ability of a randomly driven recurrent system to retain information about past inputs over time, typically quantified via linear or nonlinear memory capacity measures. Such analyses focus on expressivity and information retention properties of untrained dynamical systems, rather than on the optimization dynamics of recurrent weights under gradient-based learning. By contrast, the present work addresses a complementary question: how training dynamics themselves impose horizon-dependent limits on learnability through the interaction between temporal Jacobian transport, effective learning rates, and noise. In particular, our framework does not rely on fixed reservoirs, but instead characterizes how credit assignment and detectability degrade under heavy-tailed gradient noise, limiting the realizable horizon even in architectures with sufficient representational capacity.
@@
In the broader learning systems literature, trainability has been studied via mean-field theory, spectral initialization, curvature, and the geometry of parameter landscapes [47, 44, 35, 11, 40, 36, 48, 56, 57, 4, 28]. In recurrent and transformer settings, recent work further links trainability to Jacobian spectra, Fisher information, and anisotropic learning rates [55, 2, 50, 14]. Related analyses study gradient signal-to-noise ratios and noise scales as a function of batch size, curvature, and optimizer dynamics, with the aim of understanding optimization efficiency and generalization behavior [41, 25]. Zhou et al. [58] investigate the generalization gap between stochastic gradient descent (SGD) and Adam, attributing Adam’s performance to a sharpness reduction effect caused by adaptive coordinate-wise scaling and noting an inherent clipping behavior. However, these approaches typically analyze state dynamics and parameter dynamics in isolation. They do not treat learning as a coupled dynamical system in which temporal transport through recurrent Jacobians interacts with stochastic optimization noise, nor do they characterize the resulting universal scaling regimes governing temporal learnability and the sample complexity of detecting long-range dependencies.
@@
Simsekli et al. [49] show that mini-batch SGD noise often exhibits $\alpha$-stable, heavy-tailed behavior rather than Gaussian tails. Subsequent research has further examined the algorithmic and theoretical consequences of such heavy-tailed gradient statistics. Hübler et al. [22] establish a unified framework connecting gradient clipping and normalization, proving that both act as implicit variance control mechanisms for $\alpha$-stable noise and can stabilize training without sacrificing convergence speed. Finally, Liu and He [37] extend these ideas to online convex optimization with heavy-tailed stochastic gradients, providing convergence guarantees and adaptive algorithms that remain stable even when gradient moments of order two do not exist. In our work, we adopt the empirical reality of heavy-tailed gradient noise as a basic modeling premise.
@@ 4 Generalized effective learning rates under adaptive optimization
5 Learnability under heavy-tailed gradient noise
[…]
@@ 5.1 The envelope
bypassing the first-order expansion. This would yield $H$ numbers with no structural decomposition into gating and mixing contributions, making it impossible to assign a time scale to individual neurons. The first-order expansion provides this interpretability without losing numerical quality: as verified in Appendix C, the envelope $f(\ell)$ in Eq. 8 preserves the decay profile of $f_{\mathcal{M}}(\ell)$ (Spearman $\rho\geq 0.979$, Pearson $r\geq 0.82$, mean $0.96$).
@@ 5.3 Binary detection and matched statistic
Because the same realization of $w$ is used for all models and training runs, differences in detectability across architectures reflect differences in temporal credit assignment and gradient transport, rather than artifacts of the projection itself. Moreover, since $w$ is drawn uniformly from the unit sphere, its distribution is rotationally invariant. As a result, the projection $\langle B^{(q)}_{\ell}(\theta),\,w\rangle$ provides an unbiased random probe of the typical magnitude and relative alignment of the neuronwise parameter-sensitivity directions, without privileging any fixed parameter direction. More generally, random projections preserve inner-product and norm structure in high dimension with high probability (the Johnson–Lindenstrauss phenomenon [27]), so the signal-to-noise ratio of the projected statistic faithfully reflects that of the full gradient.
@@
Framed this way, the question of whether RNNs can use BPTT to exploit a dependency at lag $\ell$ becomes a question of detectability: is the signal induced by the state at lag $\ell$ large enough, relative to gradient noise, to remain statistically distinguishable? This viewpoint allows us to apply information-theoretic tools to characterize when past states remain recoverable during training.
@@
The factor $\mathrm{sgn}(m_{q}(\ell))$ flips each coordinate so that its contribution is oriented along the direction of its expected alignment. Consequently, under the data distribution, $\mathrm{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}$ has nonnegative mean, and every term in the sum contributes constructively. [^2] Formally, $\mathbb{E}\big[\mathrm{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}\big]=|m_{q}(\ell)|$ by definition of $m_{q}(\ell)$. Thus the expected signal contained in $S_{t,\ell}$ is always nonnegative, irrespective of the sign pattern of the raw alignments $m_{q}(\ell)$. The sign $\mathrm{sgn}(m_{q}(\ell))$ should be understood as a population-level orientation defining an ideal matched statistic; in practice, detectability depends only on the magnitude of the mean shift and is assessed from the empirical average of $S_{t,\ell}$, without requiring prior knowledge of this sign.
@@
Averaging $S_{t,\ell}$ over $N$ independent training sequences yields $\widehat{S}_{N}(\ell)=\tfrac{1}{N}\sum_{n=1}^{N}S^{(n)}_{t,\ell}$, the empirical matched statistic that forms the basis for the finite-sample analysis in the next section.

### 28 Mar 2026 v8

@@ Abstract
We develop a statistical theory of temporal learnability in recurrent neural networks, quantifying the maximal temporal horizon $\mathcal{H}_{N}$ over which gradient-based learning can recover lag-dependent structure at finite sample size $N$. The theory is built on the effective learning rate envelope $f(\ell)$, a functional that captures how gating mechanisms and adaptive optimizers jointly shape the coupling between state-space transport and parameter updates during Backpropagation Through Time. Under heavy-tailed ($\alpha$-stable) fluctuations, where empirical averages concentrate at rate $N^{-1/\kappa_{\alpha}}$ with $\kappa_{\alpha}=\alpha/(\alpha-1)$, the interplay between envelope decay and statistical concentration yields explicit scaling laws for the growth of $\mathcal{H}_{N}$: logarithmic, polynomial, and exponential temporal learning regimes emerge according to the decay law of $f(\ell)$. These results identify the envelope decay as the key determinant of temporal learnability: slower attenuation of $f(\ell)$ enlarges the learnability window $\mathcal{H}_{N}$, while heavy-tailed noise compresses temporal horizons by weakening statistical concentration. Experiments across multiple gated architectures and optimizers corroborate these structural predictions.
@@ 1 Introduction
Recurrent neural networks (RNNs) are fundamental models for processing sequential data, yet their ability to learn long-range temporal dependencies remains only partially understood. Gated architectures such as the LSTM and GRU have dramatically improved numerical stability and empirical performance, but it is still unclear which temporal dependencies are statistically recoverable under finite data. Existing analyses have studied dynamical stability, spectral properties of Jacobian products, mean-field approximations, and forward signal propagation through deep networks. These contributions characterize important necessary conditions—when gradients do not explode or vanish, or when input information survives layer-to-layer—but they typically treat state dynamics and parameter dynamics in isolation. None provides a unified statistical criterion for when transported gradient signals remain distinguishable from noise during training, which requires analyzing the coupled evolution of recurrent states and optimizer-driven parameter updates.
@@
Our previous work showed that gating induces heterogeneous time scales that shape both state evolution and gradient transport [37]. Here we develop a statistical theory of finite-horizon learnability that analyzes this coupled system. The central objects are the effective learning rates $\mu_{t,\ell}$, which quantify how Backpropagation Through Time (BPTT) reweights gradient signals across temporal lags for each neuron. These quantities arise from the interaction between gate-induced transport of gradients through recurrent dynamics and the base learning rates imposed by the optimizer. Their aggregate magnitude defines the envelope $f(\ell)=\|\mu_{t,\ell}\|_{1}$, which governs the effective strength of lagged gradient contributions to parameter updates during training. Although effective learning rates were originally [37] derived under SGD with a fixed global learning rate $\mu$, we generalize them here to adaptive optimizers such as Adam, where each parameter receives its own time-varying rate. The generalization replaces the global $\mu$ with a neuron-specific adaptive base rate $\Lambda^{(q)}_{r,\ell}$, obtained by projecting the parameter-space preconditioner onto each neuron via a Rayleigh-quotient construction, while leaving the gate-induced transport factors unchanged.
@@
The central question addressed in this paper is: given a finite number of training sequences, up to which temporal horizon can dependencies be statistically detected? To answer this, we construct a matched statistic that aggregates gradient-derived contributions across neurons and time steps, and introduce the learnability window $\mathcal{H}_{N}$, defined as the largest lag for which this statistic remains distinguishable from its noise floor at sample size $N$. Under heavy-tailed ($\alpha$-stable) fluctuations of the matched statistic, empirical averages concentrate at rate $N^{-1/\kappa_{\alpha}}$ with $\kappa_{\alpha}=\alpha/(\alpha-1)$. We show that the interaction between this concentration rate and the decay geometry of the envelope $f(\ell)$ determines the scaling behavior of $\mathcal{H}_{N}$. In particular, exponential, polynomial, and logarithmic attenuation of $f(\ell)$ induce qualitatively distinct growth laws for the learnability window, providing a universal classification of temporal learning regimes.
@@
Although the detectability relation analytically separates two factors, the decay geometry of the envelope and the statistical concentration of the matched statistic, these mechanisms are not independent during training. The temporal transport of gradients is shaped jointly by gating dynamics and optimizer adaptation, while the fluctuation statistics of the matched statistic evolve with the optimization process. Empirically, regimes exhibiting slow envelope decay often co-occur with persistent heavy-tailed fluctuations, reflecting the coupled evolution of state and parameter dynamics.
@@
Importantly, heavy-tailed noise does not enhance learnability. Instead, it weakens statistical concentration and therefore raises the data requirements for reliable detection of long-range dependencies. This interplay is reflected in the relation $N(\ell)\propto f(\ell)^{-\kappa_{\alpha}}$, which links envelope decay to the sample complexity of temporal credit assignment. Temporal learnability is thus determined by the balance between envelope attenuation and noise concentration: heterogeneous time-scale structure tends to slow the decay of $f(\ell)$ and extend the learnability window, whereas heavy-tailed fluctuations counteract this effect by reducing statistical efficiency.
@@
Relative to prior dynamical analyses of gated RNNs, this paper makes four contributions. (i) We formalize the learnability window $\mathcal{H}_{N}$ as a finite-sample measure of recoverable temporal dependencies and incorporate heavy-tailed ($\alpha$-stable) fluctuations of the matched statistic into the analysis of recurrent training. (ii) We derive explicit scaling laws that classify temporal learning regimes into canonical classes, linking envelope decay $f(\ell)$, sample-complexity requirements $N(\ell)$, and the growth of the learnability window $\mathcal{H}_{N}$. (iii) We generalize the effective learning rate framework from SGD to adaptive optimizers by introducing a neuron–lag projection of the parameter-space preconditioner. (iv) We empirically validate these structural predictions across multiple gated architectures.
@@ 2 Related work
The introduction of gating mechanisms in LSTM [20] and GRU [9] was pivotal for controlling temporal credit assignment in recurrent neural networks. Foundational and later theoretical works linked the forget gate to interpretable exponential decay and controllable memory retention [15, 50], while more recent studies place recurrent dynamics in continuous-time or implicit-model perspectives [5, 18]. These contributions illuminate how gates shape state dynamics, but do not analyze how gating multiplicative structures affect parameter updates and the capacity for learning long-range dependencies. Our framework bridges that gap by deriving per-lag effective learning rates from products of Jacobians and linking them to finite-horizon learnability.
@@
The classical vanishing/exploding gradients problem motivated many stabilization techniques: clipping, spectral regularization, and orthogonality or unitary RNNs [3, 43, 1, 53, 26]. Hierarchical, dilated, or skip architectures shorten propagation paths [29, 6, 10], while continuous-time and implicit recurrent approaches offer alternative stability–expressivity tradeoffs [5, 18]. Works on dynamical isometry control the conditioning of gradient transport [7, 44] but do not address whether transported gradients still carry statistically reliable information over long lags. Recent work further argues that vanishing and exploding gradients are not the whole story: as memory increases, parameter perturbations can induce increasingly large output variations, making gradient-based optimization highly sensitive even without gradient explosion, while element-wise recurrences and careful parametrizations help mitigate this effect [60]. In contrast, we analyze statistical detectability, showing that even when products of Jacobians remain numerically stable, learning can be ineffective if the envelope decays too rapidly. Related difficulties have long been observed in tasks with long-delayed supervision [3], where the learning signal must be propagated across long temporal gaps; however, existing analyses typically focus on optimization stability rather than statistical detectability across lags.
@@
A related but distinct line of work concerns memory capacity in reservoir computing and echo state networks, where the recurrent dynamics are kept fixed and only a readout layer is trained [24, 13, 52]. In this setting, classical results characterize the ability of a randomly driven recurrent system to retain information about past inputs over time, typically quantified via linear or nonlinear memory capacity measures. Such analyses focus on expressivity and information retention properties of untrained dynamical systems, rather than on the optimization dynamics of recurrent weights under gradient-based learning. By contrast, the present work addresses a complementary question: how training dynamics themselves impose horizon-dependent limits on learnability through the interaction between temporal Jacobian transport, effective learning rates, and noise. In particular, our framework does not rely on fixed reservoirs, but instead characterizes how credit assignment and detectability degrade under heavy-tailed fluctuations, limiting the realizable horizon even in architectures with sufficient representational capacity.
@@
In the broader learning systems literature, trainability has been studied via mean-field theory, spectral initialization, curvature, and the geometry of parameter landscapes [46, 44, 34, 11, 40, 35, 47, 56, 57, 4, 27]. In recurrent and transformer settings, recent work further links trainability to Jacobian spectra, Fisher information, and anisotropic learning rates [55, 2, 49, 14]. A complementary line of work studies stochastic optimization itself as a dynamical system. Mandt et al. [39] interpret constant-step SGD as approximate Bayesian inference with an Ornstein–Uhlenbeck-type stationary regime, while Yaida [54] establishes fluctuation–dissipation relations governing stationary SGD statistics. More recent work further analyzes the stationary distributions induced by stochastic optimization and the balance between noise and drift in high-dimensional systems [59]. Related analyses study gradient signal-to-noise ratios and noise scales as a function of batch size, curvature, and optimizer dynamics, with the aim of understanding optimization efficiency and generalization behavior [41, 25]. Zhou et al. [58] analyze the SGD–Adam generalization gap through a Lévy-driven SDE perspective, arguing that Adam’s coordinate-wise adaptation and exponential averaging can slow escape from sharp basins. However, these approaches typically analyze state dynamics and parameter dynamics in isolation. They do not treat learning as a coupled dynamical system in which temporal transport through recurrent Jacobians interacts with stochastic optimization noise, nor do they characterize the resulting universal scaling regimes governing temporal learnability and the sample complexity of detecting long-range dependencies.
@@
Simsekli et al. [48] show that mini-batch SGD noise often exhibits $\alpha$-stable, heavy-tailed behavior rather than Gaussian tails. Subsequent research has further examined the algorithmic and theoretical consequences of such heavy-tailed gradient statistics. Hübler et al. [22] connect gradient clipping to normalized SGD under heavy-tailed noise and derive improved convergence guarantees. Finally, Liu [36] studies online convex optimization with heavy tails, proving regret guarantees for classical methods under finite $p$-th moments. In our work, we adopt the empirical reality of heavy-tailed gradient noise as a basic modeling premise.
@@ 4 Generalized effective learning rates under adaptive optimization
5 Learnability under heavy-tailed fluctuations
In this section, we quantify finite-horizon learnability. Starting from the per-lag, per-neuron generalized effective learning rates $\mu_{t,\ell}^{(q)}$, we cast lag-$\ell$ credit assignment as a binary detection problem on a matched statistic built from BPTT. We then model the fluctuations of this statistic with a symmetric $\alpha$-stable ($\mathcal{S}\alpha\mathcal{S}$) location family, justify information-theoretic lower bounds via local asymptotic normality for $\alpha>1$, and translate them into sample complexity requirements and a learnability window $\mathcal{H}_{N}$. Finally, we derive scaling laws for $\mathcal{H}_{N}$ under logarithmic, polynomial, and exponential decay of the envelope $f(\ell)$, highlighting the role of the tail index $\alpha$. [^1] Throughout, we suppress the explicit dependence on the time index $t$ in quantities such as $f(\ell)$ and $m_{q}(\ell)$. In the theoretical development, $t$ is fixed and plays no role beyond anchoring the starting point of the Jacobian product. In the empirical evaluation, we average $\mu_{t,\ell}$ over all valid $t$ and all diagnostic sequences, so that only the lag dependence $\ell$ remains. In both cases, the learnability analysis depends solely on how these functionals vary with $\ell$, not on their absolute position along the sequence.
@@ 5.1 The envelope
bypassing the first-order expansion. This would yield $H$ numbers with no structural decomposition into gating and mixing contributions, making it impossible to assign a time scale to individual neurons. The first-order expansion provides this interpretability without losing numerical quality: as verified in Appendix C, the envelope $f(\ell)$ in Eq. 8 preserves the decay profile of $f_{\mathcal{M}}(\ell)$ (Spearman $\rho\geq 0.972$, Pearson $r\geq 0.823$, mean $0.95$).
@@ 5.3 Binary detection and matched statistic
Because the same realization of $w$ is used for all models and training runs, differences in detectability across architectures reflect differences in temporal credit assignment and gradient transport, rather than artifacts of the projection itself. Moreover, since $w$ is drawn uniformly from the unit sphere, its distribution is rotationally invariant. As a result, the projection $\langle B^{(q)}_{\ell}(\theta),\,w\rangle$ provides an unbiased random probe of the typical magnitude and relative alignment of the neuronwise parameter-sensitivity directions, without privileging any fixed parameter direction. Using the same realization of $w$ across models and training runs therefore yields a consistent comparative diagnostic of norm and alignment structure, without requiring that a single scalar projection exactly preserve the signal-to-noise ratio of the full gradient.
@@
Framed this way, the question of whether RNNs can use BPTT to exploit a dependency at lag $\ell$ becomes a question of detectability: is the signal induced by the state at lag $\ell$ large enough, relative to the underlying stochastic fluctuations, to remain statistically distinguishable? This viewpoint allows us to apply information-theoretic tools to characterize when past states remain recoverable during training.
@@
The factor $\operatorname{sgn}(m_{q}(\ell))$ reorients the $q$-th coordinate according to the sign of its population mean alignment, removing cancellations arising from the sign pattern of $m_{q}(\ell)$. [^2] Indeed, $\mathbb{E}\!\left[\operatorname{sgn}(m_{q}(\ell))\,\zeta^{(q)}_{t,\ell}\right]=|m_{q}(\ell)|\;\geq 0$ since $m_{q}(\ell)=\mathbb{E}[\zeta^{(q)}_{t,\ell}$]. Note that the full summand in (13) also involves $\mu^{(q)}_{t,\ell}$, which can be signed; the reorientation eliminates cancellations due to $m_{q}(\ell)$ but not those due to $\mu^{(q)}_{t,\ell}$. This sign-oriented construction should be interpreted as the theoretical matched statistic. In the empirical pipeline below, the population signs $\operatorname{sgn}(m_{q}(\ell))$ are not available, so we instead work with the raw signed sum and assess detectability through the magnitude of its empirical mean shift. A parallel theory could also be formulated for the raw signed statistic, whose expectation would still factorize into the same envelope term multiplied by a signed alignment coefficient. We introduce the sign-oriented version as an oracle benchmark because the reorientation removes cancellations due solely to heterogeneous sign patterns in the unitwise expected alignments, thereby isolating a cleaner notion of lag-$\ell$ signal strength.
@@
For finite-sample detection, the independent statistical units are full training sequences rather than individual time positions within a sequence. We therefore assign one normalized lag-specific observation to each sequence by averaging $S_{t,\ell}$ over the admissible positions $t=\ell+1,\dots,T$. This removes the trivial dependence on the number $T$-$\ell$ of valid lag positions and yields a sequence-level statistic that is comparable across lags. For the $n$th sequence, define
$\bar{S}^{(n)}_{\ell}=\frac{1}{T-\ell}\sum_{t=\ell+1}^{T}S^{(n)}_{t,\ell}$. (16)
Averaging these sequence-level statistics over $N$ independent training sequences yields
$\widehat{S}_{N}(\ell)=\tfrac{1}{N}\sum_{n=1}^{N}\bar{S}^{(n)}_{\ell},$ (17)
the finite-sample matched statistic associated with the theoretical construction above.